OpenStudy (anonymous):

I will medal you. The speed of light is 1.86 × 10^5 miles per second. If the distance between the Sun and Earth is 9.3 × 10^7 miles, how long does it take sunlight to reach Earth? 500 seconds 2 seconds 0.002 seconds 50 seconds

OpenStudy (igreen):

\((1.86 \times 10^5) \div (9.3 \times 10^7)\)

OpenStudy (igreen):

Simplify: \((9.3 \times 10^7) \div (1.86 \times 10^5)\) (Sorry this is what you have to divide, I messed up) \(93,000,000 \div 186,000\) Can you divide that? @Kyliejones